# An Empirical Study of Conformal Prediction in LLM with ASP Scaffolds for Robust Reasoning

### v1

@@ 5.5 In context learning
We designed two ICL prompts, each containing two exemplar pairs from the StepGame dataset. These prompts are comprehensively detailed in Appendix B. The first ICL prompt includes examples with 2 and 4 reasoning steps respectively, representing intermediate levels of reasoning complexity. The second ICL prompt consists of two 1-hop examples, focusing on single-step reasoning tasks. The selection of prompts is directly aligned with the calibration sets: the prompt containing 2 and 4-hop examples is utilised when calibrating with the mixed set (1-5 hops), while the prompt with 1 reasoning step is employed when calibrating with the single-length set. This distinction is repeated in the baseline experiments too, to give a fairer comparison to the CLM results.
@@ 6.1 Experiment 1
The baseline experiments were instantiated using the same LLM, where we provided the same ICL prompts as for the CLM sampling but then only sampled once, greedily, maintaining all other hyperparameters. The LLM-output is post-processed and then fed to clingo to generate answer sets. We consider a sample program correct iff there is only a single answer set generated by clingo and it contains only one answer-predicate and that answer predicate is the correct one.
@@
One notable aspect of the experiment is the evaluation on the 15-hop test set, which falls outside the range of both calibration settings. Here, the performance drops significantly and the generated answer sets are either null or mostly incorrect. This result underscores the inherent challenge of generalizing to significantly longer sequences without specific calibration data. It suggests that, while CLM with diverse calibration improves generalisation, there are still considerable difficulties in handling sequences that require substantially more reasoning steps than those seen during calibration. Finally, the improvements in performance with CLM underscore the capability of ASP to effectively capture and reason through complex logical relationships.
@@ 7 Conclusions
In this study, we explored the application of CLM to enhance the performance of a mid-range open-weight language model on complex reasoning tasks, specifically utilising the StepGame dataset. The CLM integration significantly improved accuracy on the StepGame task. When compared to baseline models that relied on standard sampling methods, CLM consistently outperformed across all tested segments. We also observe that the LLM-as-Judge metric further improved the performance of CLM, in particular on multi-step reasoning. We observed that the diversity of samples is an important consideration for the calibration set. However, we also observed that calibrating CLM on a diverse set of examples (1 to 5 reasoning steps) did not enhance generalisability to tasks requiring significantly longer reasoning steps (e.g. 15-hop). These results underscore the potential of CLM in augmenting the reasoning capabilities of standard language models, particularly when complemented by advanced evaluation metrics like LLM-as-Judge. However, the diminished performance on higher-complexity tasks shows the potential limitations of CLM. Future work could focus on a thorough examination of the statistics related to conformal set sizes, such as their distribution and relationship with task complexity. This could inform more effective sampling strategies and constraint settings within the CLM framework. Repeating the experiments multiple times to enable the calculation of error bars would also be desirable. Finally, a further baseline could be provided consisting of the LLM alone, without recourse to the external ASP reasoner.

### v2

@@ 5.5 In context learning
We designed two ICL prompts, each containing two exemplar pairs from the StepGame dataset. These prompts are comprehensively detailed in Appendix B. The first ICL prompt includes examples with 2 and 4 reasoning steps respectively, representing intermediate levels of reasoning complexity. The second ICL prompt consists of two 1-hop examples, focusing on single-step reasoning tasks. The selection of prompts is directly aligned with the calibration sets: the prompt containing 2 and 4-hop examples is used when calibrating with the mixed set (1-5 hops), while the prompt with 1 reasoning step is employed when calibrating with the single-length set. This distinction is repeated in the baseline experiments too, to give a fairer comparison to the CLM results.
@@ 6.1 Experiment 1
The baseline experiments were instantiated using the same LLM, where we provided the same ICL prompts as for the CLM sampling but then only sampled once, greedily, maintaining all other hyperparameters. The LLM-output is post-processed and then fed to clingo to generate answer sets. Note that we consider a sample program correct iff there is only a single answer set generated by clingo and it contains only one answer-predicate and that answer predicate is the correct one.
@@
One notable aspect of the experiment is the evaluation on the 15-hop test set, which falls outside the range of both calibration settings. Here, the performance drops significantly and the generated answer sets are either null or mostly incorrect. This result underscores the inherent challenge of generalising to significantly longer sequences without specific calibration data. It suggests that, while CLM with diverse calibration improves generalisation, there are still considerable difficulties in handling sequences that require substantially more reasoning steps than those seen during calibration. Finally, the improvements in performance with CLM underscore the capability of ASP to effectively capture and reason through complex logical relationships.
@@ 7 Conclusions
In this study, we explored the application of CLM to enhance the performance of a mid-range open-weight language model on complex reasoning tasks, specifically using the StepGame dataset. The CLM integration significantly improved accuracy on the StepGame task. When compared to baseline models that relied on standard sampling methods, CLM consistently outperformed across all tested segments. We also observe that the LLM-as-Judge metric further improved the performance of CLM, in particular on multi-step reasoning. We observed that the diversity of samples is an important consideration for the calibration set. However, we also observed that calibrating CLM on a diverse set of examples (1 to 5 reasoning steps) did not enhance generalisability to tasks requiring significantly more reasoning steps (e.g. 15-hop). These results underscore the potential of CLM in augmenting the reasoning capabilities of standard language models, particularly when complemented by advanced evaluation metrics like LLM-as-Judge. However, the diminished performance on higher-complexity tasks shows the potential limitations of CLM. Future work could focus on a thorough examination of the statistics related to conformal set sizes, such as their distribution and relationship with task complexity. This could inform more effective sampling strategies and constraint settings within the CLM framework. Repeating the experiments multiple times to enable the calculation of error bars would also be desirable. Finally, a further baseline could be provided consisting of the LLM alone, without recourse to the external ASP reasoner.